# A Unified Linear Programming Framework for Offline Reward Learning from Human Demonstrations and Feedback

### v1

@@ 3.2 Optimality Guarantee for Offline IRL
$D_{B}:=\left\{d\in\mathbb{R}_{+}^{|S||A|}\mid Md=(1-\gamma)\mu_{0},\;\|w_{d}\|_{\infty}\leq B\right\}.$ (3.12)
@@
$d_{\text{min}}:=\min_{(s,a)\in S\times A:\;d_{e}(s,a)\neq 0}d_{e}(s,a),$ (3.13)
since $w_{d}(s,a)\leq d_{\text{min}}^{-1}$ for any $d\in\Delta(S\times A)$ and $(s,a)\in S\times A$ by the definition of $w_{d}$ (3.3). In this case, optimality over the set $D_{B}$ implies global optimality.
It is worth highlighting that setting $B=d_{\text{min}}^{-1}$ yields results comparable to those in recent works [23, 25]. The error bounds in these works depend on the constant $\pi_{\text{min}}^{-1}$, defined as
$\pi_{\text{min}}:=\min_{(s,a)\in S\times A:\;\pi_{e}(a|s)\neq 0}\pi_{e}(a|s).$ (3.14)
@@
$d_{\text{min}}=d_{e}(s^{\prime},a^{\prime})=\pi_{e}(a^{\prime}|s^{\prime})\sum_{a\in A}d_{e}(s^{\prime},a)\geq(1-\gamma)\mu_{0}(s^{\prime})\pi_{\text{min}},$ (3.15)
@@
Our goal is to establish the optimality of $d_{e}$ within the confidence set $D_{B}$. The proof comprises two distinct steps. Firstly, we establish that $w_{\tilde{d}}$ is feasible to the dual empirical LP (3.8) with high probability for any $\tilde{d}\in D_{B}$, under appropriate level of relaxation $\epsilon_{x}$. Next, we show that $w_{e}$ has a (nearly) higher objective than $w_{\tilde{d}}$ with high probability since $w_{e}$ has a small duality gap. In the following lemma, we prove that for any $\tilde{d}\in D_{B}$, corresponding $w_{\tilde{d}}$ is a feasible solution to the empirical LP with high probability under appropriate relaxation level $\epsilon_{x}$ in the constraint.
Lemma 2.
In dual empirical LP (3.8), let
$\epsilon_{x}=\Bigg{(}B(1+\gamma)\gamma^{H}+B(1+\gamma)(1-\gamma^{H})\sqrt{\frac{2|S||A|}{N}\log\frac{2N_{x}}{\delta}}\Bigg{)}\mathbf{1},$ (3.16)
where $\delta>0$. Then, for any $\tilde{d}\in D_{B}$, $w_{\tilde{d}}$ is feasible to (3.8) with probability at least $1-\frac{\delta}{2^{|S||A|}}$.
@@
From the above Lemma, we establish the optimality guarantee in the following theorem. Specifically, in words, under the reward function $r$ recovered from the set (3.11), we show that $d_{e}$ is an $\tilde{O}(\sqrt{|S||A|/N})$-suboptimal solution over the confidence set $D_{B}$ with high probability.
Theorem 1.
Suppose $(u_{\mathcal{D}},v_{\mathcal{D}})\in\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g})$, with the relaxation level $\epsilon_{x}$ specified in Lemma 2. Let $r$ satisfy
$r(s,a)=\frac{u_{\mathcal{D}}(s,a)}{\hat{d}_{e}(s,a)}$ (3.17)
@@
$\mathbb{P}(r^{\top}d_{e}\geq r^{\top}\tilde{d}-\epsilon,\;\;\forall\tilde{d}\in D_{B})\geq 1-3\delta,$ (3.18)
@@
$\epsilon=\epsilon_{g}+(1+B)\gamma^{H}+(1-\gamma^{H})\sqrt{\frac{2}{N}\log\frac{1}{\delta}}+B(1-\gamma^{H})\sqrt{\frac{2|S||A|}{N}\log\frac{2}{\delta}}.$ (3.19)
@@
The proposed solution set achieves the $\tilde{O}(B(1-\gamma^{H})\sqrt{|S||A|/N})$ sample complexity bound with additional error terms $\epsilon_{g}$ and $(1+B)\gamma^{H}$. Note that the parameter $\epsilon_{g}$ can be set to $\epsilon_{g}=\tilde{O}(1/\sqrt{N})$ to match the sample complexity. The term $(1+B)\gamma^{H}$ diminishes exponentially with the horizon of the collected trajectory data, which underscores the requirement for long-horizon data to ensure accurate estimation. To the best of our knowledge, besides our work, [18] is the only other study that offers an optimality guarantee for the offline IRL problem under a discounted MDP, with a sample complexity of $\tilde{O}(1/\sqrt{N})$. However, as their algorithm is based on a bi-level optimization approach and their error bound is given for the log-likelihood function, a direct comparison of their result with ours is not feasible. The concurrent work by [25] presents a comparable sample complexity for an episodic MDP. We provide a detailed comparison in Appendix D.
@@
In our formulation, there exists a trade-off between the optimality of the policy and the feasibility of the reward function, which is modulated by the parameter $\epsilon_{g}$. The duality gap bound, denoted as $\epsilon_{g}$, adjusts the size of the solution set ^$\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g})$; this set expands with an increase in $\epsilon_{g}$. $\epsilon_{g}$ can be reduced to 00 without causing infeasibility, as the set ^$\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g})$ is always non-empty due to the trivial solution $(u,v)=(0,0)$. A smaller value of $\epsilon_{g}$ enhances the optimality of the expert policy $\pi_{e}$, as stated in Theorem 1. However, excessively reducing $\epsilon_{g}$ can lead to overly greedy choices, resulting in trivial or degenerate solutions. The impact of varying $\epsilon_{g}$ is demonstrated through numerical experiments in Section 5.
@@
The proposed LP formulation is well-suited for function approximation (parameterization), which allows us to reduce both the computational cost and the size (dimension) of the solution set. Consider the parameterization of the variable $(u,v)$ as $(u_{\theta},v_{\theta})$, where $\theta$ represents a parameter within the parameter space $\Theta\subset\mathbb{R}^{k}$, which we aim to explore. If there exists a $\theta\in\Theta$ such that $(u_{\theta},v_{\theta})\in\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g})$, then the optimality guarantee provided in Theorem 1 is preserved for the reward function recovered from $u_{\theta}$, while the computational complexity of the LP can be reduced to a polynomial in $k$, down from $|S||A|$. It is important to note that this formulation remains a linear (or convex) program under linear (or convex) parameterization. If non-convex function approximation is employed for high-dimensional or continuous state-action spaces, an efficient algorithm for solving the proposed optimization may be required; however, such an extension is beyond the scope of this work, and we defer this to future research.
@@ 3.3 Degeneracy Issue in Reward Learning
In the practical applications of reward learning, estimating the feasible reward set is not enough; we need to select a single reward function within the estimated feasible reward set to use. This is not a trivial problem due to existence of degenerate reward functions in the feasible reward set. Degenerate reward functions (e.g. $r=\mathbf{0}$), though theoretically feasible, are practically undesirable as they fail to separate the expert policy $\pi_{e}$ from others. In our solution set ^$\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g})$ (3.11), degeneracy in the feasibility constraint $K_{\mathcal{D}}^{\top}Xv\geq u$ is critical. If equality holds for some state-action pairs such that $(K_{\mathcal{D}}^{\top}Xv-u)(s,a)=0$, then the complementary slackness condition will not be violated by changing the value of $w_{e}(s,a)$, meaning that $w_{e}$ may not be uniquely optimal. We suggest a simple and tractable method to obtain a non-degenerate reward function in ^$\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g})$.
@@
$\begin{split}\max_{r,u,v}\quad&r^{\top}(\hat{d}_{e}-\hat{d}_{\text{sub}})\\ \text{s.t.}\quad&(u,v)\in\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g}),\quad u=\hat{d}_{e}\circ r.\end{split}$ (3.20)
@@ 4.1 LP Formulation of Offline RLHF
$\mathcal{L}(\tau^{n,1},\tau^{n,2},y^{n};r):=r^{\top}(\psi^{n,2}-\psi^{n,1})\mathbf{1}_{\{y^{n}=1\}}+r^{\top}(\psi^{n,1}-\psi^{n,2})\mathbf{1}_{\{y^{n}=2\}}.$ (4.2)
@@
^$\hat{\mathcal{R}}_{\text{HF}}(\epsilon_{r}):=\{r\mid\mathcal{L}(\tau^{n,1},\tau^{n,2},y^{n};r)\leq\epsilon_{r}\;\forall n=1,2,\ldots,N_{q},\quad r\in[-1,1]^{|S||A|}\}.$ (4.3)
@@ 4.2 Generalization Guarantee for Offline RLHF
Recent works in offline RLHF, such as [19, 20], have proposed a pessimistic MLE algorithm and provided an error bound between the estimated and the true reward function of the supposed preference model. Our LP method does not offer an optimality guarantee in the same way as these works, as it obtains a set of reward functions without assuming a specific preference model. Instead, we analyze the generalization property of the obtained reward functions by examining how $r\in\hat{\mathcal{R}}_{\text{HF}}(\epsilon_{r})$ aligns with unseen trajectory pairs sampled from $\mu_{\text{HF}}$.
We first introduce the probabilistic preference model for a human evaluator generating feedback data. We emphasize that our proposed method is not dependent on this model; we introduce it to analyze a generalization property. Suppose that $y\in\{1,2\}$ is sampled from a Bernoulli distribution with the probabilistic model $\mathbb{P}(y=1\mid\tau^{1},\tau^{2})=\Phi(r_{\text{true}}^{\top}(\psi^{1}-\psi^{2}))$, where $\Phi:\mathbb{R}\mapsto[0,1]$ is a monotonically non-decreasing function satisfying $\Phi(x)+\Phi(-x)=1$ for all $x\in\mathbb{R}$. $\Phi$ represents the preference model of the evaluator, based on their personal reward function $r_{\text{true}}$. For example, if $\Phi$ is a sigmoid function, i.e. $\Phi(x)=1/(1+e^{-x})$, then the above probabilistic model is reduced to the Bradley-Terry-Luce (BTL) model [2]. In the following theorem, we provide a generalization guarantee of any reward functions $r$ contained in the estimated solution set ^$\hat{\mathcal{R}}_{\text{HF}}(\epsilon_{r})$. Specifically, for a random (unseen) trajectory pair $(\tau^{1},\tau^{2})$ sampled from the sampling distribution $\mu_{\text{HF}}$ and the human feedback $y$ sampled from the preference model $\Phi$, we prove that the error $\mathcal{L}(\tau^{1},\tau^{2},y;r)$ is bounded by $\epsilon_{r}$ with high probability.
Theorem 2.
@@
$\mathbb{P}\left(\mathcal{L}(\tau^{1},\tau^{2},y;r)\geq\epsilon_{r}\right)\leq\sqrt{\frac{1}{2N_{q}}\log\frac{1}{\delta}}$ (4.4)
@@ 4.3 Integration of IRL and RLHF
^$\hat{\mathcal{R}}_{\text{IRL-HF}}(\epsilon_{g},\epsilon_{r})=\{(r,u,v)\mid(u,v)\in\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g}),\;r\in\hat{\mathcal{R}}_{\text{HF}}(\epsilon_{r}),\;u=\hat{d}_{e}\circ r\}.$ (4.5)
@@
If we can impose a strict reward gap between two trajectories, then the RLHF constraint can mitigate degeneracy more effectively. For instance, the constraint $r(\tau^{1})\geq r(\tau^{2})+\delta$ eliminates degenerate solutions that satisfy $r(\tau^{1})=r(\tau^{2})$ from the solution set, if we can set a strict reward gap $\delta>0$ based on human feedback data. To formalize this, we extend our approach to include the continuous feedback case, wherein the feedback is given as a continuous variable, $y\in[-1,1]$, instead of the discrete variable used in previous sections. In Appendix G, we specify this continuous feedback setting, and suggest how we can extend the definition of the solution set ^$\hat{\mathcal{R}}_{\text{HF}}(\epsilon_{r})$ as well as the generalization guarantee provided in Theorem 2. Additionally, in the next section, we compare the effects of discrete and continuous human feedback through numerical experiments.

### v3

@@ 3.2 Optimality Guarantee for Offline IRL
$D_{B}:=\left\{d\in\mathbb{R}_{+}^{|S||A|}\mid Md=(1-\gamma)\mu_{0},\;\|w_{d}\|_{\infty}\leq B\right\}.$ (18)
@@
$d_{\text{min}}:=\min_{(s,a)\in S\times A:\;d_{e}(s,a)\neq 0}d_{e}(s,a),$ (19)
since $w_{d}(s,a)\leq d_{\text{min}}^{-1}$ for any $d\in\Delta(S\times A)$ and $(s,a)\in S\times A$ by the definition of $w_{d}$ (9). In this case, optimality over the set $D_{B}$ implies global optimality.
It is worth highlighting that setting $B=d_{\text{min}}^{-1}$ yields results comparable to those in recent works (Metelli et al., 2023; Zhao et al., 2023). The error bounds in these works depend on the constant $\pi_{\text{min}}^{-1}$, defined as
$\pi_{\text{min}}:=\min_{(s,a)\in S\times A:\;\pi_{e}(a|s)\neq 0}\pi_{e}(a|s).$ (20)
@@
$\begin{split}d_{\text{min}}&=d_{e}(s^{\prime},a^{\prime})=\pi_{e}(a^{\prime}|s^{\prime})\sum_{a\in A}d_{e}(s^{\prime},a)\\ &\geq(1-\gamma)\mu_{0}(s^{\prime})\pi_{\text{min}},\end{split}$ (21)
@@
Our goal is to establish the optimality of $d_{e}$ within the confidence set $D_{B}$. The proof comprises two distinct steps. Firstly, we establish that $w_{\tilde{d}}$ is feasible to the dual empirical LP (14) with high probability for any $\tilde{d}\in D_{B}$, under appropriate level of relaxation $\epsilon_{x}$. Next, we show that $w_{e}$ has a (nearly) higher objective than $w_{\tilde{d}}$ with high probability since $w_{e}$ has a small duality gap. In the following lemma, we prove that for any $\tilde{d}\in D_{B}$, corresponding $w_{\tilde{d}}$ is a feasible solution to the empirical LP with high probability under appropriate relaxation level $\epsilon_{x}$ in the constraint.
Lemma 3.2.
In dual empirical LP (14), let
$\begin{split}\epsilon_{x}=&\Bigg{(}B(1+\gamma)\gamma^{H}+B(1+\gamma)(1-\gamma^{H})\\ &\times\sqrt{\frac{2|S||A|}{N}\log\frac{2N_{x}}{\delta}}\Bigg{)}\mathbf{1},\end{split}$ (22)
where $\delta>0$. Then, for any $\tilde{d}\in D_{B}$, $w_{\tilde{d}}$ is feasible to (14) with probability at least $1-\frac{\delta}{2^{|S||A|}}$.
@@
From the above Lemma, we establish the optimality guarantee in the following theorem. Specifically, in words, under the reward function $r$ recovered from the set (17), we show that $d_{e}$ is an $\tilde{O}(\sqrt{|S||A|/N})$-suboptimal solution over the confidence set $D_{B}$ with high probability.
Theorem 3.3.
Suppose $(u_{\mathcal{D}},v_{\mathcal{D}})\in\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g})$, with the relaxation level $\epsilon_{x}$ specified in Lemma 3.2. Let $r$ satisfy
$r(s,a)=\frac{u_{\mathcal{D}}(s,a)}{\hat{d}_{e}(s,a)}$ (23)
@@
$\mathbb{P}(r^{\top}d_{e}\geq r^{\top}\tilde{d}-\epsilon,\;\;\forall\tilde{d}\in D_{B})\geq 1-3\delta,$ (24)
@@
$\begin{split}\epsilon=\epsilon_{g}+(1+B)\gamma^{H}+(1-\gamma^{H})\sqrt{\frac{2}{N}\log\frac{1}{\delta}}\\ +B(1-\gamma^{H})\sqrt{\frac{2|S||A|}{N}\log\frac{2}{\delta}}.\end{split}$ (25)
@@
The proposed solution set achieves the $\tilde{O}(B(1-\gamma^{H})\sqrt{|S||A|/N})$ sample complexity bound with additional error terms $\epsilon_{g}$ and $(1+B)\gamma^{H}$. Note that the parameter $\epsilon_{g}$ can be set to $\epsilon_{g}=\tilde{O}(1/\sqrt{N})$ to match the sample complexity. The term $(1+B)\gamma^{H}$ diminishes exponentially with the horizon of the collected trajectory data, which underscores the requirement for long-horizon data to ensure accurate estimation. To the best of our knowledge, besides our work, (Zeng et al., 2023) is the only other study that offers an optimality guarantee for the offline IRL problem under a discounted MDP, with a sample complexity of $\tilde{O}(1/\sqrt{N})$. However, as their algorithm is based on a bi-level optimization approach and their error bound is given for the log-likelihood function, a direct comparison of their result with ours is not feasible. The concurrent work by (Zhao et al., 2023) presents a comparable sample complexity for an episodic MDP. We provide a detailed comparison in Appendix D.
@@
In our formulation, there exists a trade-off between the optimality of the policy and the feasibility of the reward function, which is modulated by the parameter $\epsilon_{g}$. The duality gap bound, denoted as $\epsilon_{g}$, adjusts the size of the solution set ^$\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g})$; this set expands with an increase in $\epsilon_{g}$. $\epsilon_{g}$ can be reduced to 00 without causing infeasibility, as the set ^$\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g})$ is always non-empty due to the trivial solution $(u,v)=(0,0)$. A smaller value of $\epsilon_{g}$ enhances the optimality of the expert policy $\pi_{e}$, as stated in Theorem 3.3. However, excessively reducing $\epsilon_{g}$ can lead to overly greedy choices, resulting in trivial or degenerate solutions. The impact of varying $\epsilon_{g}$ is demonstrated through numerical experiments in Section 5.
@@
The proposed LP formulation is well-suited for function approximation (parameterization), which allows us to reduce both the computational cost and the size (dimension) of the solution set. Consider the parameterization of the variable $(u,v)$ as $(u_{\theta},v_{\theta})$, where $\theta$ represents a parameter within the parameter space $\Theta\subset\mathbb{R}^{k}$, which we aim to explore. If there exists a $\theta\in\Theta$ such that $(u_{\theta},v_{\theta})\in\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g})$, then the optimality guarantee provided in Theorem 3.3 is preserved for the reward function recovered from $u_{\theta}$, while the computational complexity of the LP can be reduced to a polynomial in $k$, down from $|S||A|$. It is important to note that this formulation remains a linear (or convex) program under linear (or convex) parameterization. If non-convex function approximation is employed for high-dimensional or continuous state-action spaces, an efficient algorithm for solving the proposed optimization may be required; however, such an extension is beyond the scope of this work, and we defer this to future research.
@@ 3.3 Degeneracy Issue in Reward Learning
In the practical applications of reward learning, estimating the feasible reward set is not enough; we need to select a single reward function within the estimated feasible reward set to use. This is not a trivial problem due to existence of degenerate reward functions in the feasible reward set. Degenerate reward functions (e.g. $r=\mathbf{0}$), though theoretically feasible, are practically undesirable as they fail to separate the expert policy $\pi_{e}$ from others. In our solution set ^$\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g})$ (17), degeneracy in the feasibility constraint $K_{\mathcal{D}}^{\top}Xv\geq u$ is critical. If equality holds for some state-action pairs such that $(K_{\mathcal{D}}^{\top}Xv-u)(s,a)=0$, then the complementary slackness condition will not be violated by changing the value of $w_{e}(s,a)$, meaning that $w_{e}$ may not be uniquely optimal. We suggest a simple and tractable method to obtain a non-degenerate reward function in ^$\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g})$.
@@
$\begin{split}\max_{r,u,v}\quad&r^{\top}(\hat{d}_{e}-\hat{d}_{\text{sub}})\\ \text{s.t.}\quad&(u,v)\in\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g}),\quad u=\hat{d}_{e}\circ r.\end{split}$ (26)
@@ 4.1 LP Formulation of Offline RLHF
$\begin{split}\mathcal{L}(\tau^{n,1},\tau^{n,2},y^{n};r):=r^{\top}(\psi^{n,2}-\psi^{n,1})\mathbf{1}_{\{y^{n}=1\}}\\ +r^{\top}(\psi^{n,1}-\psi^{n,2})\mathbf{1}_{\{y^{n}=2\}}.\end{split}$ (28)
@@
^$\begin{split}\hat{\mathcal{R}}_{\text{HF}}(\epsilon_{r}):=\{r\mid&\mathcal{L}(\tau^{n,1},\tau^{n,2},y^{n};r)\leq\epsilon_{r}\\ &\forall n=1,2,\ldots,N_{q},\quad r\in[-1,1]^{|S||A|}\}.\end{split}$ (29)
@@ 4.2 Generalization Guarantee for Offline RLHF
Recent works in offline RLHF, such as (Zhu et al., 2023; Zhan et al., 2023), have proposed a pessimistic MLE algorithm and provided an error bound between the estimated and the true reward function of the supposed preference model. Our LP method does not offer an optimality guarantee in the same way as these works, as it obtains a set of reward functions without assuming a specific preference model. Instead, we analyze the generalization property of the obtained reward functions by examining how $r\in\hat{\mathcal{R}}_{\text{HF}}(\epsilon_{r})$ aligns with unseen trajectory pairs sampled from $\mu_{\text{HF}}$.
We first introduce the probabilistic preference model for a human evaluator generating feedback data. We emphasize that our proposed method is not dependent on this model; we introduce it to analyze a generalization property. Suppose that $y\in\{1,2\}$ is sampled from a Bernoulli distribution with the probabilistic model $\mathbb{P}(y=1\mid\tau^{1},\tau^{2})=\Phi(r_{\text{true}}^{\top}(\psi^{1}-\psi^{2}))$, where $\Phi:\mathbb{R}\mapsto[0,1]$ is a monotonically non-decreasing function satisfying $\Phi(x)+\Phi(-x)=1$ for all $x\in\mathbb{R}$. $\Phi$ represents the preference model of the evaluator, based on their personal reward function $r_{\text{true}}$. For example, if $\Phi$ is a sigmoid function, i.e. $\Phi(x)=1/(1+e^{-x})$, then the above probabilistic model is reduced to the Bradley-Terry-Luce (BTL) model (Christiano et al., 2017). In the following theorem, we provide a generalization guarantee of any reward functions $r$ contained in the estimated solution set ^$\hat{\mathcal{R}}_{\text{HF}}(\epsilon_{r})$. Specifically, for a random (unseen) trajectory pair $(\tau^{1},\tau^{2})$ sampled from the sampling distribution $\mu_{\text{HF}}$ and the human feedback $y$ sampled from the preference model $\Phi$, we prove that the error $\mathcal{L}(\tau^{1},\tau^{2},y;r)$ is bounded by $\epsilon_{r}$ with high probability.
Theorem 4.1.
@@
$\mathbb{P}\left(\mathcal{L}(\tau^{1},\tau^{2},y;r)\geq\epsilon_{r}\right)\leq\sqrt{\frac{1}{2N_{q}}\log\frac{1}{\delta}}$ (30)
@@ 4.3 Integration of IRL and RLHF
^$\begin{split}&\hat{\mathcal{R}}_{\text{IRL-HF}}(\epsilon_{g},\epsilon_{r})=\\ &\{(r,u,v)\mid(u,v)\in\hat{\mathcal{R}}_{\text{IRL}}(\epsilon_{g}),\;r\in\hat{\mathcal{R}}_{\text{HF}}(\epsilon_{r}),\;u=\hat{d}_{e}\circ r\}.\end{split}$ (31)
@@
If we can impose a strict reward gap between two trajectories, then the RLHF constraint can mitigate degeneracy more effectively. For instance, the constraint $r(\tau^{1})\geq r(\tau^{2})+\delta$ eliminates degenerate solutions that satisfy $r(\tau^{1})=r(\tau^{2})$ from the solution set, if we can set a strict reward gap $\delta>0$ based on human feedback data. To formalize this, we extend our approach to include the continuous feedback case, wherein the feedback is given as a continuous variable, $y\in[-1,1]$, instead of the discrete variable used in previous sections. In Appendix G, we specify this continuous feedback setting, and suggest how we can extend the definition of the solution set ^$\hat{\mathcal{R}}_{\text{HF}}(\epsilon_{r})$ as well as the generalization guarantee provided in Theorem 4.1. Additionally, in the next section, we compare the effects of discrete and continuous human feedback through numerical experiments.